OpenStudy (anonymous):

What is the solution set for |–2x + 16| – 9 = –3

OpenStudy (ledah):

First, isolate the absolute value operation: |–2x + 16| = 6 Simplify: |8 - x| = 3 Now, break this into the two cases of an absolute value operation: CASE 1: 8 - x >= 0, so we can simply remove the absolute value operation: 8 - x = 3 x = 8 - 3 = 5 CASE 2: 8 - x < 0, and we have to flip the sign to remove the absolute value operation: x - 8 = 3 x = 3 + 8 = 11 Each of these answers fulfills the case restriction, and satisfies the original problem. x = 5 OR x = 11 Hope this helps.

OpenStudy (anonymous):